OpenStudy (anonymous):

Solve two times square root of x plus 13 end root plus 5 equals 17

OpenStudy (anonymous):

\[2\sqrt{x+13}+\sqrt{5}=17\] \[(2\sqrt{x+13} +\sqrt{5})^2=(17)^2\] \[4(x+13)+5=17\] \[4x+52+5=17\] \[4x+57=17\] \[4x=-40 \therefore x= -10\]

OpenStudy (anonymous):

@helpme1.2 what do you think of that?

OpenStudy (anonymous):